OpenStudy (lukecrayonz):

A painter working at 100 feet from the ground drops his paint brush onto the street below. The function, s(t)=-16t^2+100, gives the height in feet of the paint brush as it falls to the ground. The velocity at time t=2 seconds is given by lim t->2 s(2)-s(t)/(2-t). Evaluate this limit to find the velocity of the paint brush at 2 seconds.

OpenStudy (lukecrayonz):

http://screensnapr.com/v/qU586F.png

OpenStudy (lukecrayonz):

Just to make clear what the equation is.

OpenStudy (lukecrayonz):

@satellite73

OpenStudy (lukecrayonz):

@amistre64 my hero?

OpenStudy (lukecrayonz):

@saifoo.khan @Hero anyone D:

OpenStudy (saifoo.khan):

dont we have to differentiate the distance equation to get the equation for velocity?

OpenStudy (lukecrayonz):

I feel like I just called up the justice league of the smartest people ever at math..

OpenStudy (amistre64):

i think its asking for first principles

OpenStudy (anonymous):

yes we could cheat and take the derivative and replace t by 2, but lukecrayons hasn't gotten there yet in calc, so we have to grind it out

OpenStudy (saifoo.khan):

Ah. okays..

OpenStudy (saifoo.khan):

What level of math is that? @Lukecrayonz

OpenStudy (lukecrayonz):

Precalculus, the final chapter, Introduction to calculus.

OpenStudy (anonymous):

-64

OpenStudy (saifoo.khan):

We can do it the easier way if we are doing calc. ;) But you aren't. :l

OpenStudy (anonymous):

\[s(t)=-16t^2+100\] \[s(2)=36\] \[s(t)-s(2)=-16t^2+100-36=-16t^2+64\] \[=-16(t^2-4)=-16(t+2)(t-2)\]

OpenStudy (saifoo.khan):

lol^

OpenStudy (anonymous):

-64 ft/s

OpenStudy (anonymous):

so \[\frac{s(2)-s(t)}{2-t}=\frac{16(2-t)(2+t)}{2-t}\]\] \[=16(2+t)\] take the limit as t goes to 2, get \[16(2+2)=16\times 4=64\]

OpenStudy (anonymous):

something like that should be negative though, because it is falling

OpenStudy (anonymous):

yes it's negative

OpenStudy (anonymous):

you forgot a sign -

OpenStudy (saifoo.khan):

Hey is latex working? D:

sam (.sam.):

@saifoo.khan yes

OpenStudy (anonymous):

never stopped for me must be you!

OpenStudy (lukecrayonz):

Thank you all!

OpenStudy (saifoo.khan):

This is what i am getting from a long time.

OpenStudy (saifoo.khan):

I will try to reload.

OpenStudy (lukecrayonz):

Since you guys are here.. I have another question..

OpenStudy (lukecrayonz):

http://screensnapr.com/v/caYPLG.png

OpenStudy (lukecrayonz):

actually I'll repost.